# Gelation in Vector Multiplicative Coalescence and Extinction in Multi-Type Poisson Branching Processes

### v1

@@ 3. Correspondence between Gelation and Extinction Probability with Applications
when $s_{i}=y_{i}/(\alpha_{i}t)$.
Using this equivalence we get our first application that yields the gelation phenomenon for the vector multiplicative coalescent process and the new proof follows immediately from this equivalence, Theorem 2.1 and equation (12).
@@
(a) If $\rho(VD[\alpha_{i}]t)\leq 1$, then the total mass $\sum_{\bf x}\zeta_{\bf x}(t)|{\bf x}\rangle=\boldsymbol{\alpha}$.
(b) If $\rho(VD[\alpha_{i}]t)>1$, then the total mass $\sum_{\bf x}\zeta_{\bf x}(t)|{\bf x}\rangle<\boldsymbol{\alpha}$.
@@
Remark 3.3.
Let us note that an additional relationship between vector multiplicative coalescence and multi-type branching processes through Burgers’ equation was investigated in a recent paper [10].

### v2

@@ 5. Correspondence between Gelation and Extinction Probability with Applications
when $s_{i}=y_{i}/(\alpha_{i}t)$. The identification of the matrix of means ${\bf M}$ in the multi-type Poisson branching process in terms of the time parameter $t$ of the vector multiplicative coalescent process implies that $t$ is the parameter for a continuous sequence of branching processes rather than the usual time (generations) view of branching processes.
Using this equivalence we get our first application that yields the gelation phenomenon for the vector multiplicative coalescent process discussed in Section 3. We state the result again in the theorem below followed by the short new proof that follows immediately from this equivalence.
@@
(a) If $\rho(VD[\alpha_{i}]t)\leq 1$, then the total mass $\sum_{\bf x}\zeta_{\bf x}(t)|{\bf x}\rangle=|{\boldsymbol{\alpha}}\rangle$.
(b) If $\rho(VD[\alpha_{i}]t)>1$, then the total mass $\sum_{\bf x}\zeta_{\bf x}(t)|{\bf x}\rangle<|{\boldsymbol{\alpha}}\rangle$.
Consequently, the gelation time for the vector multiplicative coalescent process is given by $T_{gel}={1\over\rho(VD[\alpha_{i}])}$.
Proof.
Applying the identification ${\bf M}=VD[\alpha_{i}]t$ to Theorem 4.1 yields
(a) If $\rho(VD[\alpha_{i}]t)\leq 1$, then ${\boldsymbol{\eta}}={\bf 1}$, the $k$-dimensional vector $(1,1,\ldots,1)$.
(b) If $\rho(VD[\alpha_{i}]t)>1$, then ${\boldsymbol{\eta}}$ is the smallest solution less than ${\bf 1}$ of the fixed point equations
$e^{\langle{\bf e}_{i}\,|VD[\alpha_{i}]t\,|{\bf s}-\boldsymbol{1}\rangle}=s_{i}$
Then the equivalence (19) implies that $\boldsymbol{\alpha}t\boldsymbol{\eta}={\bf y}$ which completes the proof. ∎
Remark 5.2.
In a recent work [12], Hoogendijk et al. described an alternative approach, which relies on establishing the equivalence between the Smoluchowski coagulation equations (see equation (2)) and the inviscid Burgers equation using branching process techniques.
@@
In particular, the homogenous multi-type Poisson branching process where $\boldsymbol{\alpha}={\bf 1}$ and $V$ is matrix with all entries $1$, the extinction probabilities have the form
(21) $\eta_{\ell}=\sum_{{\bf x}}\frac{1}{{\bf x}!\ {\bf x}^{\bf 1}}\ \tau\left(K_{k},x_{i}x_{j}t\right)e^{-\langle{\bf x}|\boldsymbol{1}\rangle kt}x_{\ell}\,,\,\,\,\,\,\,\forall\ell=1,2\ldots,k.$